# Randomly punctured Reed–Solomon codes achieve list-decoding capacity over linear-sized fields

### v5

@@ 2.4 Reduced intersection matrices: full column rank
The following theorem shows that reduced intersection matrices of $k$-weakly-partition-connected hypergraphs are nonsingular when viewed as a matrix over $\mathbb{F}_{q}(X_{1},\dots,X_{n})$. This was essentially conjectured by Shangguan and Tamo [ST20] and essentially established by Brakensiek, Gopi, and Makam [BGM23], who conjectured and showed, respectively, nonsingularity of the (non-reduced) intersection matrix under similar conditions. By the same union bound argument as in [ST20, Theorem 5.8], Theorem 2.11 already implies list-decodability of Reed–Solomon codes up to the generalized Singleton bound over exponentially large fields sizes, which is [BGM23, Theorem 1.5]. For completeness, and to demonstrate how the hypergraph perspective more directly connects the list-decoding problem to the GM-MDS theorem, we include a proof of Theorem 2.11 in Appendix A.
Theorem 2.11 (Full column rank. Implicit from Theorem A.2 of [BGM23]).
@@
We note that, [BGM23] assumes throughout their paper that the alphabet size $q$ is sufficiently large, but Theorem 2.11 follows from the weaker “$q$ sufficiently large” version: For any fixed field size $q$, take $Q$ to be a sufficiently large power of $q$. Then, by the “$q$ sufficiently large” version of Theorem 2.11, matrix $\mathsf{RIM}_{Q,\mathcal{H}}$ has full column rank over the field $\mathbb{F}_{Q}(X_{1},\dots,X_{n})$. Hence, the determinant of some square full-rank submatrix of $\mathsf{RIM}_{Q,\mathcal{H}}$ is a nonzero polynomial in $\mathbb{F}_{Q}[X_{1},\dots,X_{n}]$. The entries of $\mathsf{RIM}_{Q,\mathcal{H}}$ can all be viewed as polynomials over $\mathbb{F}_{q}$, so the corresponding full-rank submatrix of $\mathsf{RIM}_{q,\mathcal{H}}$ has a determinant that is a nonzero polynomial in $\mathbb{F}_{q}[X_{1},\dots,X_{n}]$ — symbolically, the determinants are the same polynomials, as $\mathbb{F}_{q}$ and $\mathbb{F}_{Q}$ have the same characteristic. Hence, the matrix $\mathsf{RIM}_{q,\mathcal{H}}$ has full column rank over the field $\mathbb{F}_{q}(X_{1},\dots,X_{n})$.
@@ 2.5 Reduced intersection matrix: row deletions
As in [GZ23], we consider row deletions from the reduced intersection matrix. The goal of this section is to establish Lemma 2.14, that the full-column-rank-ness of reduced intersection matrices are robust to row deletions.
@@
The next lemma appears in a weaker form in [GZ23]. It roughly says that, given a reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}$ with some constant factor “slack” in the combinatorial constraints, we can omit a constant fraction of the rows without compromising the full-column-rank-ness of the matrix. Our version of this lemma saves roughly a factor of $t\sim L$ compared to the analogous lemma [GZ23, Lemma 3.11]. The reason is that the $k$-weakly-partition-connected condition is more robust to these row deletions (by a factor of roughly $t$) than the condition in [GZ23]. As such, our proof is also more direct.
Lemma 2.14 (Robustness to deletions. Similar to Lemma 3.11 of [GZ23]).
Let $\mathcal{H}=([t],\mathcal{E})$ be a $(k+\varepsilon n)$-weakly-partition-connected hypergraph with $t\geq 2$. For all sets $B\subset[n]$ with $|B|\leq\varepsilon n$, we have that $\mathsf{RIM}_{\mathcal{H}}^{B}$ is nonempty and has full column rank.
@@
$\sum_{e\in\mathcal{E}^{\prime}}{\max\{|\mathcal{P}(e)|-1,0\}}$ $=\sum_{i\in[n]}{\max\{|\mathcal{P}(e)|-1,0\}}-\sum_{i\in B}{\max\{|\mathcal{P}(e)|-1,0\}}$
@@ 3.1 Overview of the proof
$e_{n-2}$$e_{n-1}$$e_{n}$$e_{n-2}$$e_{n-1}$$e_{n}$ Lemma 2.8 Bad list-decoding configuration has $(k+\varepsilon n)$-w.p.c agreement hypergraph Lemma 2.3 RIMs for agreement hypergraphs do not have full column rank Lemma 3.1 RIMs for $(k+\varepsilon n)$-w.p.c hypergraphs have full column rank w.h.p. Theorem 1.1 RS code list-decodable w.h.p. ptUnion bound over possibleagreement hypergraphs Lemma 3.8 If RIM not full column rank, it admits a certificate. Corollary 3.10 Number of possible certificates is small. Corollary 3.12 The probability of any one certificate is very small ptUnion bound overpossible certificates
Properties of GetCertificate, which generates certificates for non-full-rank RIMs.
@@
By Lemma 2.8 and Lemma 2.3, every bad list-decoding configuration admits a weakly-partition-connected agreement hypergraph whose reduced intersection matrix does not have full column rank. Thus, to prove Theorem 1.1, it suffices to show that, with high probability, every such reduced intersection matrix has full column rank. The main technical lemma for this section is the one stated below. Our main result, Theorem 1.1, follows by applying Lemma 2.3 and Lemma 2.8 with Lemma 3.1, and taking a union bound over all $\sum_{t=2}^{L+1}{2^{tn}}$ possible agreement hypergraphs.
@@
At the highest level, the proof of Lemma 3.1 follows the same outline as [GZ23]. For every sequence of evaluation points $(\alpha_{1},\dots,\alpha_{n})\in\mathbb{F}_{q}^{n}$ for which $\mathsf{RIM}_{\mathcal{H}}$ does not have full column rank, we show that there is a certificate $(i_{1},\dots,i_{r})\in[n]^{r}$ consisting of distinct indices in $[n]$ (Lemma 3.8), which intuitively “attests” to the failure of the matrix $\mathsf{RIM}_{\mathcal{H}}$ to be full column rank. We then show that, for any certificate $(i_{1},\dots,i_{r})$, the probability that $(\alpha_{1},\dots,\alpha_{n})$ has certificate $(i_{1},\dots,i_{r})$ is exponentially small. (More precisely, it will at most be $(\frac{(t-1)k}{q-n})^{r}$. See Corollary 3.12). We then show that there are not too many certificates (Corollary 3.10), and then union bound over the number of possible certificates to obtain the desired result (Lemma 3.1).
Our argument differs from [GZ23] in how we choose our certificates. The argument of [GZ23] allowed for up to $n^{r}$ certificates. Our argument instead only needs $\binom{n}{r}2^{tr}$ many certificates, which is much smaller when $r=\Omega(n)$ (the parameter regime of interest here) and overall allows us to save a factor of $n$ in the alphabet size. Our savings comes from leveraging that there are at most $2^{t}$ different “types” of hyperedges (see Remark 2.9), and thus at most $2^{t}$ different types of variables $X_{i}$ in the reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}$. This observation was alluded to in [GZ23]. [^8] Guo and Zhang [GZ23] write “It is possible that achieving an alphabet size linear in n would require establishing and exploiting other properties of intersection matrices or reduced intersection matrices, such as an appropriate notion of exchangeability.” We found this prediction to be insightful and true. With this observation in mind, we assume, without loss of generality, that the edges of $\mathcal{H}$ are ordered by their respective type (we can relabel the edges of $\mathcal{H}$, which effectively permutes the rows of $\mathsf{RIM}_{\mathcal{H}}$).
@@
Our method of generating a certificate $(i_{1},\dots,i_{r})$ for the evaluation sequence $(\alpha_{1},\dots,\alpha_{n})$ (Algorithm 2) is similar to that of [GZ23] at a high level—with each certificate $i_{1},\dots,i_{r}$, we associate a sequence of $(t-1)k\times(t-1)k$ submatrices $M_{1},\dots,M_{r}$ of $\mathsf{RIM}_{\mathcal{H}}$ (Algorithm 1) that are entirely specified by $i_{1},\dots,i_{r}$ as follows: since evaluating $X_{[n]}=\alpha_{[n]}$ forces $\mathsf{RIM}_{\mathcal{H}}$ to not be full rank, then so will all of its $(t-1)k\times(t-1)k$ submatrices. Thus if we sequentially ’reveal’ $X_{1}=\alpha_{1},X_{2}=\alpha_{2},\dots$, then at some point, $M_{j}$ becomes singular exactly when we set $X_{i_{j}}=\alpha_{i_{j}}$ — in fact, $i_{j}$ is defined as such, so that we select $M_{1},i_{1},M_{2},i_{2},\dots,$ in that order, but we emphasize that $M_{j}$ can be computed from $i_{1},\dots,i_{j-1}$ without knowing $\alpha_{1},\dots,\alpha_{n}$. Conditioned on $M_{j}$ being non-singular with $X_{1}=\alpha_{1},\dots,X_{i_{j}-1}=\alpha_{i_{j}-1}$, the probability that $M_{j}$ becomes singular when setting $X_{i_{j}}=\alpha_{i_{j}}$ is at most $\frac{(t-1)k}{q-n}$: $\alpha_{i_{j}}$ is uniformly random over at least $q-n$ field elements, and the degree of $X_{i_{j}}$ in the determinant of $M_{j}$ is at most $(t-1)k$ (and the determinant is nonzero by definition). Running conditional probabilities in the correct order, we conclude that the probability that a particular certificate $i_{1},\dots,i_{r}$ is generated is at most $(\frac{(t-1)k}{q-n})^{r}$, just as in [GZ23].
@@
Whereas [GZ23] pick any matrix $M_{j}$ that is obtained after removing the variables $X_{i_{1}},\ldots,X_{i_{j-1}}$, we do a more deliberate choice of matrices by leveraging the symmetries of $\mathsf{RIM}_{\mathcal{H}}$ (Remark 2.9). First, we ensure that we can keep a “bank” of $\Omega_{t}(r)$ unused variables of each of the $O_{t}(1)$ types. Then, starting with a full column rank submatrix $M$ of $\mathsf{RIM}_{\mathcal{H}}$ devoid of all variables in the “bank,” we start sequentially applying the evaluations $X_{1}=\alpha_{1},X_{2}=\alpha_{2},\ldots$. Whenever $M(X_{\leq i_{1}}=\alpha_{\leq i_{1}})$ turns singular, we find that the evaluation $X_{i_{1}}=\alpha_{i_{1}}$ is what ’caused’ it to become singular. We then go to the “bank” to find a variable $X_{i_{1}^{\prime}}$ of the same type as $X_{i_{1}}$ and “re-indeterminate” $M$ by replacing all instances of $X_{i_{1}}$ in $M$ with $X_{i_{1}^{\prime}}$. That way, we ensure that $M$ is, in a sense, “reused.” Furthermore, we ensure $i_{1}^{\prime}>i_{1}$, so that the matrix $M(X_{\leq i_{1}}=\alpha_{\leq i_{1}})$ is now nonsingular, so we can keep going. Of course, if we end up reaching the end (i.e. $M(X_{[n]}=\alpha_{[n]})$ is full column rank), then in fact, $\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ is full column rank, and so the evaluations $(\alpha_{1},\ldots,\alpha_{n})$ were ‘good’ after all.
@@
Otherwise, if the evaluations $(\alpha_{1},\ldots,\alpha_{n})$ were ‘bad’, then the submatrix $M$ couldn’t have reached the end, and that can only happen if some specific type was completely exhausted from the bank. However, given the size of our initial bank, that must have meant that $M$ must have been “re-indeterminated” at least $\Omega_{t}(r)$ times. When that happens, we collect the indices $i_{1},\dots,i_{\ell}$ that we gathered from this round, remove them from $\mathsf{RIM}_{\mathcal{H}}$, and repeat the process again with a refreshed bank. Since we only need $r$ indices, then we end up doing at most $O_{t}(1)$ rounds. Because each round yields a strictly increasing sequence of indices of length at least $\Omega_{t}(r)$, then we up getting a certificate consisting of at most $O_{t}(1)$ strictly increasing runs of total length $r$, of which there are at most $\binom{n}{r}\cdot O_{t}(1)^{r}$.
@@
To be more concrete, when we generate the submatrix $M=M_{1}$, we ensure that any variable appearing in $M_{1}$ has the same type as $\Omega_{t}(r)$ variables that are not in $M_{1}$ (but still in $\mathsf{RIM}_{\mathcal{H}}$). This creates a “bank” of variables of each type. Then, if $X_{\leq i_{1}}=\alpha_{\leq i_{1}}$ was the point that made $M_{1}$ singular, we can get $M_{2}$ by replacing all copies of $X_{i_{1}}$ with some $X_{i_{1}^{\prime}}$ that is of the same type and in the “bank.” Since variables $i_{1}$ and $i_{1}^{\prime}$ are of the same type, they have analogous rows in the reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}$, so this new matrix $M_{2}$ is still a submatrix of $\mathsf{RIM}_{\mathcal{H}}$. Therefore, we can pick up where we left off with $M_{1}$ but with $M_{2}$ instead. That is, $M_{2}$ will in fact be full rank when we apply the evaluations $X_{\leq i_{1}}=\alpha_{\leq i_{1}}$. Thus the next index $i_{2}$ on which $M_{2}$ turns singular will be strictly greater than $i_{1}$. We then repeat the process in $M_{2}$, replacing $X_{i_{2}}$ with some $X_{i_{2}^{\prime}}$ that is in the “bank” and of the same type, getting $M_{3}$, and so on. We can continue this process for $\Omega_{t}(r)$ steps because of the size of the bank of each type, so we get an increasing run of length $\Omega_{t}(r)$ in our certificate. After we run out of some type in our bank, we remove the used indices $i_{1},\dots,i_{\ell}$ from $\mathsf{RIM}_{\mathcal{H}}$ and repeat the process again with a refreshed bank. This continues for $O_{t}(1)$ times only, as we only need $r$ indices in the end.
We now finish the proof of Theorem 1.1, assuming Lemma 3.1. The rest of this section is devoted to proving Lemma 3.1.
Proof of Theorem 1.1, assuming Lemma 3.1.
By Lemma 2.3, if $RS_{n,k}(\alpha_{1},\dots,\alpha_{n})$ is not $\left(\frac{L}{L+1}(1-R-\varepsilon),L\right)$ average-radius list-decodable, then there exists a vector $y$ and pairwise distinct codewords $c^{(1)},\dots,c^{(t)}$ with $t\geq 2$ such that the agreement hypergraph $\mathcal{H}=([t],\mathcal{E})$ is $(R+\varepsilon)n=(k+\varepsilon n)$-weakly-partition-connected. By Lemma 2.8, the matrix $\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ is not full column rank. Now, the number of possible agreement hypergraphs $\mathcal{H}$ is at most $\sum_{t=2}^{L+1}2^{tn}\leq 2^{(L+2)n}$. Thus by the union bound over possible agreement hypergraphs $\mathcal{H}$ with Lemma 3.1, we have, for $r=\lfloor{\frac{\varepsilon n}{2}}\rfloor$,
[…]
$\leq\mathop{\bf Pr\/}_{\alpha_{[n]}}\left[\exists\text{ $(k+\varepsilon n)$-w.p.c. agreement hypergraph }\mathcal{H}\text{ such that }\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})\text{ not full column rank}\right]$
$\leq 2^{(L+2)n}\max_{\text{$(k+\varepsilon n)$-w.p.c. }\mathcal{H}}\quad\mathop{\bf Pr\/}_{\alpha_{[n]}}\left[\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})\text{ not full column rank}\right]$
as desired. Here, we used that $q=n+k\cdot 2^{10L/\varepsilon}$. ∎
@@
We now devote the rest of this Section to proving Lemma 3.1.
3.2.0.0.1 Types.
For a hypergraph $\mathcal{H}=([t],(e_{1},\dots,e_{n}))$, the type of an index $i$ (or, by abuse of notation, the type of the variable $X_{i}$, or the edge $e_{i}$) is simply the set $e_{i}\subset[t]$. There are $2^{t}$ types, and by abuse of notation, we identify the types by the numbers $1,2,\dots,2^{t}$ in an arbitrary fixed order with a bijection $\tau:\text{(subsets of $[t]$)}\to[2^{t}]$. We say a hypergraph is type-ordered if the hyperedges $e_{1},\dots,e_{n}$ are sorted according to their type: $\tau(e_{1})\leq\tau(e_{2})\leq\cdots\leq\tau(e_{n})$. Since permuting the labels of the edges of $\mathcal{H}$ preserves the rank of $\mathsf{RIM}_{\mathcal{H}}$ (it merely permutes the rows of $\mathsf{RIM}_{\mathcal{H}}$), we can without loss of generality assume in Lemma 3.1 that $\mathcal{H}$ is type-ordered.
3.2.0.0.2 Global variables.
@@
1 $B\leftarrow\emptyset$, $i_{0}\leftarrow\perp$, $\ell_{0}\leftarrow\perp$ for $\ell=1,\dots,j$ do

### v6

@@ 2.4 Reduced Intersection Matrices: Full Column Rank
The following theorem shows that reduced intersection matrices of $k$-weakly-partition-connected hypergraphs are nonsingular when viewed as a matrix over $\mathbb{F}_{q}(X_{1},\dots,X_{n})$. This was essentially conjectured by Shangguan and Tamo [59] and essentially established by Brakensiek, Gopi, and Makam [9], who conjectured and showed, respectively, nonsingularity of the (non-reduced) intersection matrix under similar conditions. By the same union bound argument as in [59, Theorem 5.8], Theorem 2.11 already implies list-decodability of Reed–Solomon codes up to the generalized Singleton bound over exponentially large field sizes, which is [9, Theorem 1.5]. For completeness, and to demonstrate how the hypergraph perspective more directly connects the list-decoding problem to the GM-MDS theorem, we include a proof of Theorem 2.11 in Appendix A.
Theorem 2.11 (Full column rank. Implicit from Theorem A.2 of [9]).
@@
We note that, [9] assumes throughout their paper that the alphabet size $q$ is sufficiently large, but, as in Theorem 2.11, this assumption is easily dropped: For any fixed field size $q$, take $Q$ to be a sufficiently large power of $q$. Then, by the “$q$ sufficiently large” version of Theorem 2.11, matrix $\mathsf{RIM}_{Q,\mathcal{H}}$ has full column rank over the field $\mathbb{F}_{Q}(X_{1},\dots,X_{n})$. Hence, the determinant of some square full-rank submatrix of $\mathsf{RIM}_{Q,\mathcal{H}}$ is a nonzero polynomial in $\mathbb{F}_{Q}[X_{1},\dots,X_{n}]$. The entries of $\mathsf{RIM}_{Q,\mathcal{H}}$ can all be viewed as polynomials over $\mathbb{F}_{q}$, so the corresponding full-rank submatrix of $\mathsf{RIM}_{q,\mathcal{H}}$ has a determinant that is a nonzero polynomial in $\mathbb{F}_{q}[X_{1},\dots,X_{n}]$ — symbolically, the determinants are the same polynomials, as $\mathbb{F}_{q}$ and $\mathbb{F}_{Q}$ have the same characteristic. Hence, the matrix $\mathsf{RIM}_{q,\mathcal{H}}$ has full column rank over the field $\mathbb{F}_{q}(X_{1},\dots,X_{n})$.
@@ 2.5 Reduced Intersection Matrix: Row Deletions
We consider row deletions from the reduced intersection matrix. The goal of this section is to establish Lemma 2.14, that the full-column-rank-ness of reduced intersection matrices are robust to row deletions.
@@
The next lemma roughly says that, given a reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}$ with some constant factor “slack” in the combinatorial constraints, we can omit a constant fraction of the rows without compromising the full-column-rank-ness of the matrix.
Lemma 2.14 (Robustness to deletions).
Let $\mathcal{H}=([t],\mathcal{E})$ be a $(k+\varepsilon n)$-weakly-partition-connected hypergraph with $t\geq 2$, where $\mathcal{E}=(e_{1},\dots,e_{n})$. For all sets $B\subseteq[n]$ with $|B|\leq\varepsilon n$, we have that $\mathsf{RIM}_{\mathcal{H}}^{B}$ is nonempty and has full column rank.
@@
$\sum_{e\in\mathcal{E}^{\prime}}{\max\{|\mathcal{P}(e)|-1,0\}}$ $=\sum_{i\in[n]}{\max\{|\mathcal{P}(e_{i})|-1,0\}}-\sum_{i\in B}{\max\{|\mathcal{P}(e_{i})|-1,0\}}$
@@ 3.1 Overview of the Proof
Lemma 2.8 Bad list-decoding configuration has $(k+\varepsilon n)$-w.p.c agreement hypergraph Lemma 2.3 RIMs for agreement hypergraphs do not have full column rank Lemma 3.1 RIMs for $(k+\varepsilon n)$-w.p.c hypergraphs have full column rank w.h.p. Theorem 1.1 RS code list-decodable w.h.p. Union bound over possible agreement hypergraphs Lemma 3.8 If RIM not full column rank, it admits a certificate. Corollary 3.10 Number of possible certificates is small. Corollary 3.12 The probability of any one certificate is very small Union bound over possible certificates Properties of GetCertificate, which generates certificates for non-full-rank RIMs.
@@
To explain our ideas, we first consider why previous results in [59, 25, 9] require an exponentially large alphabet. By Lemma 2.3 and Lemma 2.8, every bad list-decoding configuration admits a weakly-partition-connected agreement hypergraph whose reduced intersection matrix does not have full column rank. Thus, to prove list-decodability, it suffices to show that, with high probability, every such reduced intersection matrix has full column rank. Because these matrices have full column rank as symbolic matrices (Theorem 2.11, see also [9]), they have full column rank under random evaluations with probability at least $1-\frac{Lk^{2}}{q}$ by the Schwarz-Zippel lemma (take the determinant of any full rank square submatrix). Then, if $q$ is sufficiently large, by a union bound over all $2^{(L+1)n}$ possible reduced intersection matrices, all reduced intersection matrices have full column rank with high probability, so our code is list-decodable with high probability. This union bound requires $q$ to be at least exponential in $n$.
Our key high-level insight is to use the “slackness” in the coding parameters to amplify the probability that the reduced intersection matrix fails to be full rank from $O_{L,k,n}(1/q)$ to $1/q^{\Omega(n)}$; that way, a smaller alphabet size $q$ suffices. To see this, consider the toy problem of independently picking $m$ random row vectors $v_{1},\cdots,v_{m}\in\mathbb{F}_{q}^{n}$ to form an $m\times n$ matrix $M$, which we want to have full column rank. If we choose $m=n$, which is the optimal choice of $m$, then the probability that M has full column rank is bounded by a function that is $\Theta(1/q)$, and this happens only if each $v_{i}$ is not in the span of $v_{1},\dots,v_{i-1}$. However, suppose we choose $m=(1+\varepsilon)n$ for some small $\varepsilon>0$. In this case, we could afford $\varepsilon n$ “faulty” vectors $v_{i}$ , i.e., $v_{i}$ may be in the span of previous vectors, in which case we just skip it and consider the next vector. The probability that the matrix $M$ has full column rank is then exponentially small, $1/q^{\Omega(\varepsilon n)}$. In the same way, the slackness in our coding parameters means the reduced intersection matrix has roughly $1+\varepsilon$ times as many rows as columns, so (with some additional ideas) we can similarly amplify the probability that the reduced intersection matrix fails to be full rank.
Lemma 3.1 below captures this probability amplification. Our main result, Theorem 1.1, follows by applying Lemma 2.3 and Lemma 2.8 with Lemma 3.1, and taking a union bound over all $\sum_{t=2}^{L+1}{2^{tn}}$ possible agreement hypergraphs.
@@
At the highest level, our proof of Lemma 3.1 is a union bound over “not-full-column-rank certificates.” For every sequence of evaluation points $(\alpha_{1},\dots,\alpha_{n})\in\mathbb{F}_{q}^{n}$ for which $\mathsf{RIM}_{\mathcal{H}}$ does not have full column rank, we show that there is a certificate $(i_{1},\dots,i_{r})\in[n]^{r}$ of distinct indices in $[n]$ (Lemma 3.8), which intuitively “attests” to the failure of the matrix $\mathsf{RIM}_{\mathcal{H}}$ to be full column rank. We then show that, for any certificate $(i_{1},\dots,i_{r})$, the probability that $(\alpha_{1},\dots,\alpha_{n})$ has certificate $(i_{1},\dots,i_{r})$ is exponentially small. (More precisely, it will at most be $(\frac{(t-1)k}{q-n})^{r}$. See Corollary 3.12). We then show that there are not too many certificates (Corollary 3.10), and then union bound over the number of possible certificates to obtain the desired result (Lemma 3.1).
Given an evaluation sequence $(\alpha_{1},\dots,\alpha_{n})$, we deterministically generate a certificate $(i_{1},\dots,i_{r})$ as described in Algorithm 2. Alongside, we also produce a sequence of $(t-1)k\times(t-1)k$ submatrices $M_{1},\dots,M_{r}$ of $\mathsf{RIM}_{\mathcal{H}}$, as specified in Algorithm 1, with each matrix entirely determined by the indices $i_{1},\dots,i_{r}$. The first matrix, $M_{1}$, is a canonical full-rank submatrix of $\mathsf{RIM}_{\mathcal{H}}(X_{[n]})$. More generally, for each $j$, the matrix $M_{j}$ is a full-rank submatrix of $\mathsf{RIM}_{\mathcal{H}}(X_{[n]})$ and is a deterministic function of $i_{1},\dots,i_{j-1}$. Additionally, each index $i_{j}$ depends only on $M_{j}$ and the evaluation points $\alpha_{1},\dots,\alpha_{n}$. In this way, we sequentially select $M_{1},i_{1},M_{2},i_{2},\dots,$ in that order.
@@
To explain the choice of $i_{j}$, we note that evaluating $X_{[n]}=\alpha_{[n]}$ forces $\mathsf{RIM}_{\mathcal{H}}$ to not be full rank, then all of its $(t-1)k\times(t-1)k$ submatrices are not full rank. Thus if we sequentially “reveal” $X_{1}=\alpha_{1},X_{2}=\alpha_{2},\dots$, then at some point, $M_{j}$ becomes not-full-rank. We define $i_{j}$ to be the index such that setting $X_{i_{j}}=\alpha_{i_{j}}$ makes $M_{j}$ not-full-rank. Conditioned on $M_{j}$ being full rank with $X_{1}=\alpha_{1},\dots,X_{i_{j}-1}=\alpha_{i_{j}-1}$, the probability that $M_{j}$ becomes not-full-rank when setting $X_{i_{j}}=\alpha_{i_{j}}$ is at most $\frac{(t-1)k}{q-n}$: $\alpha_{i_{j}}$ is uniformly random over at least qnq-n field elements, and the degree of $X_{i_{j}}$ in the determinant of $M_{j}$ is at most (t1)k(t-1)k (and the determinant is nonzero by definition). It is possible to run the conditional probabilities in the correct order to conclude that the probability that a particular certificate $i_{1},\dots,i_{r}$ is generated is at most $(\frac{(t-1)k}{q-n})^{r}$.
At this point, we can already obtain a quadratic alphabet size by proving a weaker version of (9). The above argument works if we choose $M_{j}$ to be any canonical, say the lexicographically smallest, full rank square submatrix of $\mathsf{RIM}_{\mathcal{H}}^{\{i_{1},\dots,i_{j-1}\}}$. By a union bound over all certificates—of which there are, naively, at most $n^{r}$—the probability that $\textsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ is not full rank is at most $n^{r}\cdot(\frac{(t-1)k}{q-n})^{r}$, which is exponentially small for $q=\Theta(n^{2})$, so we get our list-decoding result with alphabet size $q=\Theta(n^{2})$. To improve the alphabet size to quadratic, we improve the bound on the number of certificates to $\binom{n}{r}2^{tr}$, which is much smaller — $2^{\Theta(n)}$ rather than $n^{\Theta(n)}$ — when $r=\Omega(n)$, the parameter regime of interest here. Our savings in the number of certificates comes from leveraging that there are at most $2^{t}$ different “types” of hyperedges (see Remark 2.9), and thus at most $2^{t}$ different types of variables $X_{i}$ in the reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}$. With this observation in mind, we assume, without loss of generality, that the edges of $\mathcal{H}$ are ordered by their respective type (we can relabel the edges of $\mathcal{H}$, which effectively permutes the rows of $\mathsf{RIM}_{\mathcal{H}}$).
@@
To reduce the number of certificates, we make a more deliberate choice of matrices $M_{j}$ that leverages the symmetries of $\mathsf{RIM}_{\mathcal{H}}$ (Remark 2.9) and forces the certificates to have $O_{t}(1)$ increasing sequences $i_{1}<i_{2}<i_{3}<\cdots$. First, we ensure that we can keep a “bank” of $\Omega_{t}(r)$ unused variables of each of the $O_{t}(1)$ types. Then, starting with a full column rank submatrix $M$ of $\mathsf{RIM}_{\mathcal{H}}$ devoid of all variables in the “bank,” we start sequentially applying the evaluations $X_{1}=\alpha_{1},X_{2}=\alpha_{2},\ldots$. Whenever $M(X_{\leq i_{1}}=\alpha_{\leq i_{1}})$ turns singular, we find that the evaluation $X_{i_{1}}=\alpha_{i_{1}}$ is what ’caused’ it to become singular. We then go to the “bank” to find a variable $X_{i_{1}^{\prime}}$ of the same type as $X_{i_{1}}$ and “re-indeterminate” $M$ by replacing all instances of $X_{i_{1}}$ in $M$ with $X_{i_{1}^{\prime}}$. That way, we ensure that $M$ is, in a sense, “reused.” Furthermore, we ensure $i_{1}^{\prime}>i_{1}$, so that the matrix $M(X_{\leq i_{1}}=\alpha_{\leq i_{1}})$ is now nonsingular, so we can keep going. Of course, if we end up reaching the end (i.e. $M(X_{[n]}=\alpha_{[n]})$ is full column rank), then in fact, $\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ is full column rank, and so the evaluations $(\alpha_{1},\ldots,\alpha_{n})$ were ‘good’ after all.
@@
Otherwise, if the evaluations $(\alpha_{1},\ldots,\alpha_{n})$ were ‘bad’, then the submatrix $M$ couldn’t have reached the end, and that can only happen if some specific type was completely exhausted from the bank. However, given the size of our initial bank, that must have meant that $M$ must have been “re-indeterminated” at least $\Omega_{t}(r)$ times. When that happens, we collect the indices $i_{1},\dots,i_{\ell}$ that we gathered from this round, remove them from $\mathsf{RIM}_{\mathcal{H}}$, and repeat the process again with a refreshed bank. Since we only need $r$ indices, we end up doing at most $O_{t}(1)$ rounds. Because each round yields a strictly increasing sequence of indices of length at least $\Omega_{t}(r)$, then we end up getting a certificate consisting of at most $O_{t}(1)$ strictly increasing runs of total length $r$, of which there are at most $\binom{n}{r}\cdot O_{t}(1)^{r}$ by simple counting.
@@
To be more concrete, when we generate the submatrix $M=M_{1}$, we ensure that any variable appearing in $M_{1}$ has the same type as $\Omega_{t}(r)$ variables that are not in $M_{1}$ (but still in $\mathsf{RIM}_{\mathcal{H}}$). This creates a “bank” of variables of each type. Then, if $X_{\leq i_{1}}=\alpha_{\leq i_{1}}$ was the evaluation that made $M_{1}$ singular, we can get $M_{2}$ by replacing all copies of $X_{i_{1}}$ with some $X_{i_{1}^{\prime}}$ that is of the same type and in the “bank.” Since variables $i_{1}$ and $i_{1}^{\prime}$ are of the same type, they have analogous rows in the reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}$, so this new matrix $M_{2}$ is still a submatrix of $\mathsf{RIM}_{\mathcal{H}}$. Therefore, we can pick up where we left off with $M_{1}$ but with $M_{2}$ instead. That is, $M_{2}$ will in fact be full rank when we apply the evaluations $X_{\leq i_{1}}=\alpha_{\leq i_{1}}$. Thus the next index $i_{2}$ on which $M_{2}$ turns singular will be strictly greater than $i_{1}$. We then repeat the process in $M_{2}$, replacing $X_{i_{2}}$ with some $X_{i_{2}^{\prime}}$ that is in the “bank” and of the same type, getting $M_{3}$, and so on. We can continue this process for $\Omega_{t}(r)$ steps because of the size of the bank of each type, so we get an increasing run of length $\Omega_{t}(r)$ in our certificate. After we run out of some type in our bank, we remove the used indices $i_{1},\dots,i_{\ell}$ from $\mathsf{RIM}_{\mathcal{H}}$ and repeat the process again with a refreshed bank. This continues for $O_{t}(1)$ times only, as we only need $r$ indices in the end.
@@
We devote most of the remainder of this section to formally proving Lemma 3.1, and conclude with the proof of Theorem 1.1.
Types.
For a hypergraph $\mathcal{H}=([t],(e_{1},\dots,e_{n}))$, the type of an index $i$ (or, by abuse of notation, the type of the variable $X_{i}$, or the edge $e_{i}$) is simply the set $e_{i}\subseteq[t]$. There are $2^{t}$ types, and by abuse of notation, we identify the types by the numbers $1,2,\dots,2^{t}$ in an arbitrary fixed order with a bijection $\tau:2^{[t]}\to[2^{t}]$, where $2^{[t]}$ denotes the power set of $[t]$. We say a hypergraph is type-ordered if the hyperedges $e_{1},\dots,e_{n}$ are sorted according to their type: $\tau(e_{1})\leq\tau(e_{2})\leq\cdots\leq\tau(e_{n})$. Since permuting the labels of the edges of $\mathcal{H}$ preserves the rank of $\mathsf{RIM}_{\mathcal{H}}$ (it merely permutes the rows of $\mathsf{RIM}_{\mathcal{H}}$), we can without loss of generality assume in Lemma 3.1 that $\mathcal{H}$ is type-ordered.
Global variables.
@@
1 $B\leftarrow\emptyset$, $i_{0}\leftarrow\perp$, $\ell_{0}\leftarrow\perp$
2 for $\ell=1,\dots,j$ do